Heather is a big Leaf’s fan. She decides to open a shop selling Leaf’s gear. The biggest seller will be the Leaf jerseys so Heather makes her plans based on that.

A licensed Leaf jersey will cost her $75. Store rent will cost her $10,000 a month. Since she can’t be there all the time she needs to hire two staff at $3,000 per month each.

The big question is what the price of a sweater will be. If the Leafs make the playoffs and go deep, some of the sweaters will become very popular and will sell for $250 each. If the Leafs play poorly and don’t make the playoffs, sweaters will sell for $200 each.


a)  In both a worst case and best case scenario, how many sweaters does Heather have to sell each month to make a target profit of $84,000 per year?

 

 

b)  Briefly state whether this is a good business for Heather and why.

 




1
Expert's answer
2021-03-01T11:25:09-0500

A)


Optimistic scenario (Price per PC:250$):

If Quantity sweaters is x:


Quantity sweaters = (250x - 75x -10000-6000)*12=84000

250x - 75x -10000-6000=7000

175x -16000=7000

175x =23000

x ≈131,43 PC per month

Full year Quantity sweaters for profit in the amount of the 84 000$ is 1577 PC


Pessimistic scenario(Price per PC:200$):


Quantity sweaters = (200x - 75x -10000-6000)*12=84000

200x - 75x -10000-6000=7000

125x -16000=7000

125x =23000

x ≈184 PC per month

Full year Quantity sweaters for profit in the amount of the 84 000$ is 2208 PC


B) In both case shop will have the healthy business marginality (19% for pessimistic scenario and 21% for optimistic scenario)